A company gives each of its 20 employees​ (assume they were all employed continuously through 2016 and​ 2017) 12 days of vacation a year if they are employed at the end of the year. The vacation accumulates and may be taken starting January 1 of the next year. The employees work 8 hours per day. In​ 2016, they made​ $17.50 per hour and in 2017 they made​ $20 per hour. During​ 2017, they took an average of 8 days of vacation each. The​ company's policy is to record the liability existing at the end of each year at the wage rate for that year. What amount of vacation liability would be reflected on the 2016 and 2017 balance​ sheets, respectively?

A.38,400;12,800

B.38,400​;49,600

C.33,600​;12,800

D.33,600​;49,600

''D'' is the correct answer, can you show me how to calculate it please?
Answer #1

Vacation liability for 2016 =Number of employees x vacation given x hours worked in a day x wage rate

= 20*12*8*17.5

= $33,600

2017: 4 vacation (12 - 8 used) for 2016 due + 12 current year= 16

Liability = (12*20*8*20) + (4*12*8*17.5) = $49,600